The following information pertains to the standard costs and actual activity for Tyler Company for September:

Standard cost per unit
Direct materials 4 units of material A x $ 3.00 per unit
1 unit of material B x $ 4.00 per unit
Direct labor 3 hours x $ 9.00 per hour
Activity for September
Materials purchased
Material A 4,500 units x $ 3.10 per unit
Material B 1,100 units x $ 4.15 per unit
Materials used
Material A 4,150 units
Material B 1,005 units
Direct labor used 2,950 hours x $ 9.40 per hour
Production output 1,000

There were no beginning direct materials inventories.

(a) Determine the materials price and quantity variances.

Do not round final answers.

Material A Material B
Materials price variance Answer AnswerFU Answer AnswerFU
Materials quantity variance Answer AnswerFU Answer AnswerFU

(b) Determine the labor rate and efficiency variances.

Labor rate variance Answer AnswerFU
Labor efficiency variance Answer AnswerFU

Please answer all parts of the question.

PreviousSave AnswersNext
Answer #1
For material A :
Standard quantity of materials for actual production = Actual units produced * Standard quantity of material per unit = 1000 * 4 4000
Materials price variance = ( Actual price - Standard price ) * Actual quantity = ( 3.10 - 3.00 ) * 4150 415 U
Materials quantity variance = ( Actual quantity - Standard quantity ) * Standard price = ( 4150 - 4000 ) * 3.00 450 U
For material B :
Standard quantity of materials for actual production = Actual units produced * Standard quantity of material per unit = 1000 * 1 1000
Materials price variance = ( Actual price - Standard price ) * Actual quantity = ( 4.15 - 4.00 ) * 1005 151 U
Materials quantity variance = ( Actual quantity - Standard quantity ) * Standard price = ( 1005 - 1000 ) * 4.00 20 U
Material A Material B
Materials price variance 415 U 151 U
Materials quantity variance 450 U 20 U
(b)
Standard labor hours for actual production = Actual units production * Standard labor hours per unit = 1000 * 3 3000
Labor rate variance = ( Actual rate - Standard rate ) * Actual hours = ( 9.40 - 9.00 ) * 2950 1180 U
Labor efficiency variance = ( Actual hours - Standard hours ) * Standard rate = ( 2950 - 3000 ) * 9.00 (450) F
Note : In the absence of information regarding rounding off, the answers are rounded off to the nearest whole dollar.